Which property of the real numbers is illustrated by the following statement?
10 • (18 + 3) = 10 • 18 + 10 • 3

Answer by sprolden(40)

Which property of the real numbers is illustrated by the following statement?
10 • (18 + 3) = 10 • 18 + 10 • 3
Distributive Property a(b+c)=ab+ac